eaゃalto V(x-r). qXX expression is The, lx-r
Answer #1


\sqrt{(x+y)^2-4xy}

=\sqrt{x^2+y^2+2xy-4xy}

=\sqrt{x^2+y^2-2xy}

=\sqrt{(x-y)^2}\\=x-y

Option (d) is correct.